OpenStudy (anonymous):

The accompanying table shown here display 12 observations of a pair of variables (x, y). The variables x and y are positively correlated, with a correlation coefficient of r = 0.96. You estimate the slope of the least squares regression line, y = a + bx, for these data values to be b = 1.7. What is the y-intercept, a, of the least squares regression line? Round your answer to the nearest tenth. x 3 6 8 9 9 13 15 18 20 21 23 23 y 9 17 15 15 18 24 29 36 43 39 46 35

OpenStudy (anonymous):

y = a + bx b = 1.7 y = a + 1.7x

OpenStudy (anonymous):

So do i take one of the x,y values and plug them into the equation ?